# Determination of nuclear quadrupole moments of 25Mg, 87Sr, and 135,137Ba via configuration-interaction plus coupled-cluster approach

### v1

@@ II.1 Hyperfine interaction
where $I$, $J$, and $F$ represent the nuclear spin, the electron orbital, and the total angular momentum, respectively. The state $|\gamma IJFM_{F}\rangle$ is the hyperfine state constructed by coupling a nuclear eigenstate $|IM_{I}\rangle$ with an atomic eigenstate $|\gamma JM_{J}\rangle$, where $\gamma$ represents the remaining electronic quantum numbers. The first-order correction $\Delta{E_{F}^{(1)}}$ of the hyperfine interaction to the energy can be defined as:
@@
where $R_{N}=\sqrt{5/3}\langle{r^{2}}\rangle^{1/2}$ is the radius of the sphere, and $\langle{r^{2}}\rangle^{1/2}$ is the charge root-mean-square radius of the nucleus.
@@ II.2 Configuration interaction plus coupled-cluster method
In the many-electron atomic system, electron-electron correlation encompasses core-core, core-valence, and valence-valence correlations. All three types of correlation effects are of paramount importance for the precise evaluation of hyperfine interaction properties. In the present work, we adopt a hybrid approach that integrates the relativistic configuration interaction with coupled-cluster methods to comprehensively account for these three types of electron-electron correlations. Within the framework of this hybrid methodology, coupled-cluster calculation is employed to depict the core-core and core-valence correlations, while the valence-valence correlation is taken into consideration through a configuration interaction calculation. This hybrid method was originally developed by Safronova et al. Safronova et al. (2009), and subsequently, Dzuba independently conceived a similar approach Dzuba (2014). The method utilized in our work is conceptually similar to theirs. However, we have independently developed the corresponding software package for this hybrid method, therefore there remain a few differences in the handling details compared to their methods Tang (2025).
@@
where $H_{\rm DF}$ represents the Dirac-Fock Hamiltonian, and $\Sigma_{1}$ and $\Sigma_{2}$ correspond to the one-body and two-body correlation potentials, respectively. In this work, we derive the one-body and two-body correlation potentials via second-order many-body perturbation (MBPT(2)) calculations and coupled-cluster calculations.
@@
Afterwards, one can compute the transition matrix elements of various operators by utilizing the obtained wave functions. During the calculation of these transition matrix elements, we take into account the random phase approximation (RPA), core Brueckner effects, structural radiation, and normalization corrections of all orders Tang et al. (2017); Blundell et al. (1989). For the convenience of description, later we will combine the core Brueckner effects, structural radiation, and normalization corrections and refer to them as the high-order (HO) correction beyond RPA. Moreover, we include the two-particle (TP) interaction up to the second order Safronova et al. (1999); Savukov (2004). A detailed account of the method and the corresponding formulas has been provided in a recent work Tang (2025).
@@ II.3 Computation details
Table 1 shows the parameters of the Gauss basis set, along with the basis spaces utilized in the coupled-cluster (CC) and configuration interaction (CI) calculations. $N$ represents the number of the Gauss functions. $N_{\rm core}$ indicates the number of core orbitals. $N_{\rm valence}$ and $N_{\rm virtual}$ respectively denote the numbers of valence orbitals and virtual orbitals involved in the CC calculation. $N_{\rm CI}$ stands for the number of orbitals employed to build the configurations. Specifically, single-particle orbitals with an energy lower than 20000 $a.u.$ are designated as virtual orbitals, and those with an energy lower than 500 $a.u.$ are used for constructing configurations. Nevertheless, in the MBPT calculations, the summation is conducted over the entire basis set.
To obtain the uncertainty of the results, we use five distinct methods to construct the one-body and two-body correlation potentials. Method 1 (CI+MBPT) derives these potentials via second-order many-ody perturbation theory. Method 2 (CI+LCCSD) builds them through linear coupled-cluster singles and doubles calculations, and Method 3 (CI+CCSD) determines them via full coupled-cluster singles and doubles calculations, with the latter considering non-linear terms related to single and double excitations of the cluster operator compared to CI+LCCSD. In Method 2 and Method 3, only contributions from single and double excited states are considered. To compensate for the overlooked higher-order correlation effects, a rescaling parameter $\rho_{\kappa}$ is introduced to replace the one-body correlation potential $\Sigma_{1}$ with $\rho_{\kappa}\Sigma_{1}$, and fine-tuning this parameter can bring the calculated energy closer to the experimental energy, a strategy previously used in CI+MBPT calculations  Zhang et al. (2023, 2024). Method 4 (CI+LCCSDs) starts from LCCSD-obtained potentials, keeps two-body potentials constant and applies a rescaling parameter to the one-body potential. Method 5 (CI+CCSDs) does the same but starts from CCSD-obtained potentials. The rescaling parameters $\rho_{\kappa}$ for CI+LCCSDs and CI+CCSDs are listed in Table 2.
@@ III.1 Energy
From Table 3, for the Mg atom, CI+MBPT, CI+LCCSD, and CI+CCSD all yield relatively accurate results, with differences from measured values less than 200 cm-1. CI+LCCSD has the smallest difference, suggesting a cancellation between unconsidered correlation effects in CI+CCSD and non-linear terms of single and double excitations. As atomic number $Z$ increases for the three atoms, the difference between CI+MBPT results and measured values rises from 200 cm-1 to 2500 cm-1. The difference for CI+LCCSD results also increases but less significantly. CI+CCSD results are relatively stable, with differences within 250 cm-1, possibly due to more complete correlation effects. CI+LCCSDs and CI+CCSDs, the two methods with scaling parameters, have even smaller differences (controlled within 100 cm-1) from measured values. However, they rely on measured values and are useful for subsequent uncertainty evaluation.
@@ III.2 Magnetic dipole hyperfine-structure constant
$3s3p$ $^{3}\!P^{o}_{1}$ –143(1) –144.945 Lurio (1962) –141.365 Veseth (1987),–146.1 Porsev and Derevianko (2004)
$3s3p$ $^{3}\!P^{o}_{2}$ –127(1) –128.440 Lurio (1962) –126.287 Veseth (1987),–129.7 Porsev and Derevianko (2004)
@@
The last column of Table 5 shows other theoretical results. For the $3s4s$ $^{3}\!S_{1}$ state of 25Mg, there’s one available result from CI+all-order calculation He et al. (2009), equivalent to our CI+LCCSD. Their method only considered RPA, while ours includes RPA, HO, and TP corrections. Considering only RPA, our CI+LCCSD result (323.8 MHz) differs from theirs (325 MHz) by about 1 MHz. For $3s3p$ $^{3}\!P^{o}_{1,2}$ states of 25Mg, Veseth’s many-body perturbation theory results Veseth (1987) differ from ours by $1\%-2\%$. The CI+MBPT results by Beloy et al. Beloy et al. (2008) are consistent with ours, with a difference of less than $1$ MHz. The CI+MBPT results by Porsev et al. Porsev and Derevianko (2004) also match our recommended values.
For the 87Sr atom, our final recommended values are very close to the MCDF results of Lu et al. Lu et al. (2019), and the CI+MBPT results of Poserv et al. Porsev and Derevianko (2004), with a difference of less than $1$ MHz. The CI+MBPT results of Beloy et al. Beloy et al. (2008) are very close to our CI+MBPT results. We have not found that the hyperfine parameters of the $5s4d$ $^{1}\!D_{2}$ state have been calculated by other theoretical methods. In contrast, in the 137Ba atom, the agreement between the CI+MBPT results of Kozlov et al.  Kozlov and Porsev (1999) and the experimental results is not as good as that of our results. This indicates that these states of the 137Ba atom are more sensitive to electron correlation effects, and it also shows that the electron correlation effects beyond the second-order many-body perturbation theory are important. In general, the configuration interaction plus coupled-cluster method can accurately calculate the hyperfine-structure properties of alkaline-earth atom systems.
@@ III.3 Nuclear quadrupole moment
Table 6: The electric-field gradients $q$ (in a.u.) of the low-lying states in 25Mg , 87Sr, and 137Ba.
@@
Table 6 shows electric-field gradients $q$ of low-lying states of Mg, Sr, and Ba atoms. Like Table 5, it lists five methods’ calculation results, final recommended values, and uncertainties. For Mg, CI+MBPT method’s results differ from the other four by about 3.8%, larger than for hfs constants $A$; differences among the other four are within 1%. In Sr’s $5s5p$ $^{3}P^{o}_{1}$ and $5s5p$ $^{3}P^{o}_{2}$ states, CI+MBPT results differ from the other four by about 10%, while the other four differ within 2%, similar to the hfs constant $A$ results in Table 4. For Sr’s $5s4d$ $^{1}\!D_{2}$ state, the electric-field gradient’s electron-correlation trend differs from the hfs constant $A$. CI+MBPT results differ from the other four by about 4%, but for hfs constant $A$, CI+MBPT result is four times smaller. In Ba’s $6s5d$ $^{3}\!D_{1}$, $6s5d$ $^{3}\!D_{2}$, and $6s5d$ $^{3}\!D_{3}$ states, the five methods’ results differ within 3%, and CI+MBPT result differs from the final recommended value within 2%. For Ba’s $6s5d$ $^{1}\!D_{2}$ and $6s6p$ $^{1}\!P^{o}_{1}$ singlet states, CI+MBPT results differ from the final recommended values by 8.5% and 25% respectively, showing importance of correlation effects beyond MBPT(2). The other four methods’ results differ within 3%. Comparing Table 4 and Table 6, there are differences in electron correlation trends between magnetic-dipole and electric-quadrupole hyperfine interactions.
@@
By combining the electric-field gradients in Table 6 and the measured electric quadrupole hyperfine-structure constants of these low-lying states Lurio (1962); Grundevik et al. (1983); Zu Putlitz (1963); Heider and Brink (1977); Gustavsson et al. (1979); Schmelling (1974); Wijngaarden and Li (1995), we obtain the nuclear quadrupole moments of 25Mg, 87Sr, 135Ba and 137Ba, which are listed in Table 7. In Table 7, we not only present the nuclear quadrupole moments extracted from individual states, but also provide the final recommended values and their uncertainties through statistical averaging of multiple states. The accuracy of these measured hfs constants $B$ is high. Therefore, the uncertainties of the nuclear quadrupole moments we extracted are entirely due to the electric-field gradients.
For 25Mg, the nuclear quadrupole moments extracted from the $3s3p$ $^{3}\!P^{o}_{1}$ and $3s3p$ $^{3}\!P^{o}_{1}$ states are very close to the final recommended value. The differences among the three are less than 0.2%, and the uncertainty of the final recommended value is approximately 1%. For 87Sr, the differences among the nuclear quadrupole moments obtained from $5s4d~^{1}\!D_{2}$ and $5s5p~^{3}\!P_{1,2}$ states are within 1%.
@@
The comparison shows that our result is in excellent agreement with the nuclear quadrupole moments extracted from the $6p_{3/2}$ state of 137Ba+ Silverans et al. (1986); Wendt et al. (1988); Sahoo (2006b). However, there is a deviation of about 4% compared with the results extracted from the $5d_{3/2,5/2}$ states of 137Ba+ Sahoo et al. (2013); Chakraborty et al. (2025). For the $6p_{3/2}$ state, two groups independently carried out the measurement of its electric quadrupole hyperfine-structure constants Wendt et al. (1988); Villemoes et al. (1993). The measurement results they obtained are in very good agreement, and the measurement precision of both groups is better than 1%. Similarly, for the $5d_{3/2,5/2}$ states, two teams independently measured the hyperfine splittings of these two states Silverans et al. (1986); Lewty et al. (2013). The values of the hfs constants $B$ they obtained are highly consistent, and the measurement precision of the $5d_{3/2,5/2}$ states is at least an order of magnitude higher than that of the $6p_{3/2}$ state. If the theoretical framework can accurately provide the electric-field gradient, then by combining the hfs constants $B$ of the $6p_{3/2}$ or $5d_{3/2,5/2}$ states, the nuclear quadrupole moment can be extracted accurately. In the recent work by Chakraborty et al. Chakraborty et al. (2025), they used the coupled-cluster singles, doubles, and triples method to calculate the electric-field gradient of the $5d_{3/2,5/2}$ states of 137Ba+, and obtained an accurate nuclear quadrupole moment of 137Ba. The uncertainty of their result is extremely small. Our result is obtained by statistically averaging the results of five states of the 137Ba atom. Except for the $6s5d$ $^{1}\!D_{2}$ state, the difference between our calculated magnetic-dipole hyperfine-structure constant and the measured value is within 2% as shown in Table 5. Based on this, we believe that our result should be reliable. However, there is still room for improvement in our method. For example, the contribution of the triple-excitation is not currently included in our CC calculation, and the three-body interaction potential is also ignored, which may be one of the reasons for the 4% difference. However, we think that for the nuclear quadrupole moments of 135,137Ba, it still needs to be further verified by other calculations or new mearsurements. Since the situation of 135Ba is the same as that of 137Ba, we only compared 137Ba in Table 8.

### v2

# Determination of nuclear quadrupole moments for 25Mg, 87Sr, and 135,137Ba via configuration-interaction combined with a coupled-cluster approach

@@ II.1 Hyperfine interaction
where $I$, $J$, and $F$ represent the nuclear spin, total angular momentum of the electrons, and the total angular momentum, respectively. The state $|\gamma IJFM_{F}\rangle$ is the hyperfine state constructed by coupling a nuclear eigenstate $|IM_{I}\rangle$ with an atomic eigenstate $|\gamma JM_{J}\rangle$, where $\gamma$ represents the remaining electronic quantum numbers. The first-order correction $\Delta{E_{F}^{(1)}}$ of the hyperfine interaction to the energy can be defined as:
@@
where $R_{N}=\sqrt{5/3}\langle{r^{2}}\rangle^{1/2}$ is the radius of the sphere, and $\langle{r^{2}}\rangle^{1/2}$ is the charge root-mean-square radius of the nucleus. In the present work, the nucleus is modeled as a uniformly magnetized and charged sphere.
@@ II.2 Configuration interaction plus coupled-cluster method
In the many-electron atomic system, electron-electron correlation encompasses core-core, core-valence, and valence-valence correlations. All three types of correlation effects are of paramount importance for the precise evaluation of hyperfine interaction properties. In the present work, we adopt a hybrid approach that integrates the relativistic configuration interaction with coupled-cluster methods to comprehensively account for these three types of electron-electron correlations. Within the framework of this hybrid methodology, coupled-cluster calculation is employed to depict the core-core and core-valence correlations, while the valence-valence correlation is taken into consideration through a configuration interaction calculation. The relativistic configuration interaction plus a linearized version of coupled-cluster theory, called the RCI+all-order method, was first developed by Safronova et al. Safronova et al. (2009); subsequently, Dzuba independently devised a similar approach Dzuba (2014). The concept of this hybrid approach can be traced back to the RCI+MBPT combination introduced by Dzuba and co-workers Dzuba et al. (1996). The method utilized in our work is conceptually similar to theirs. However, we have independently developed the corresponding software package for this hybrid method, therefore there remain a few differences in the handling details compared to their methods Tang (2025).
@@
where $H_{\rm DF}$ represents the Dirac-Fock Hamiltonian, and $\Sigma_{1}$ and $\Sigma_{2}$ correspond to the one-body and two-body correlation potentials, respectively. In this work, we derive the one-body and two-body correlation potentials via second-order many-body perturbation (MBPT(2)) calculations and coupled-cluster calculations. Here we give only the effective equation for the valence electrons; the core electrons are incorporated into the one- and two-body correlation potentials, while the interaction equations for the many-electron system are detailed in Ref. Dzuba et al. (1996).
@@
Afterwards, one can compute the transition matrix elements of various operators by utilizing the obtained wave functions. During the calculation of these transition matrix elements, we take into account the random phase approximation (RPA), core Brueckner effects, structural radiation, and normalization corrections of all orders Tang et al. (2017); Blundell et al. (1989); Dzuba et al. (1998). For the convenience of description, later we will combine the core Brueckner effects, structural radiation, and normalization corrections and refer to them as the high-order (HO) correction beyond RPA. Moreover, we include the two-particle (TP) interaction up to the second order Dzuba et al. (1998); Safronova et al. (1999); Savukov (2004). A detailed account of the method and the corresponding formulas has been provided in a recent work Tang (2025).
@@ II.3 Computation details
Table 1 shows the parameters of the Gauss basis set, along with the basis spaces utilized in the coupled-cluster (CC) and configuration interaction (CI) calculations. $N$ represents the number of the Gauss functions. $N_{\rm core}$ indicates the number of core orbitals. $N_{\rm valence}$ and $N_{\rm virtual}$ respectively denote the numbers of valence orbitals and virtual orbitals involved in the CC calculation. $N_{\rm CI}$ stands for the number of orbitals employed to build the configurations. Specifically, single-particle orbitals with an energy lower than 20000 $a.u.$ are designated as virtual orbitals, and those with an energy lower than 500 $a.u.$ are used for constructing configurations. Nevertheless, in the MBPT calculations, the summation is conducted over the entire basis set. The sizes of the Gaussian basis set, virtual orbital space, and configurations quoted in Table 1 are already fully adequate: any further increase in basis set, virtual orbitals or configurations would shift the energies and hyperfine-structure parameters by less than one unit in the last reported digit. In short, within the present theoretical framework, the results are absolutely converged in basis set, virtual orbital, and configuration space.
To obtain the uncertainty of the results, we use five methods to construct the one-body and two-body correlation potentials. Method 1 (CI+MBPT) derives these potentials via second-order many-ody perturbation theory. Method 2 (CI+LCCSD) builds them through linear coupled-cluster singles and doubles calculations, and Method 3 (CI+CCSD) determines them via full coupled-cluster singles and doubles calculations, with the latter considering non-linear terms related to single and double excitations of the cluster operator compared to CI+LCCSD. In Method 2 and Method 3, only contributions from single and double excited states are considered. To compensate for the overlooked higher-order correlation effects, a rescaling parameter $\rho_{\kappa}$ is introduced to replace the one-body correlation potential $\Sigma_{1}$ with $\rho_{\kappa}\Sigma_{1}$, and fine-tuning this parameter can bring the calculated energy closer to the experimental energy, a strategy previously used in CI+MBPT calculations  Zhang et al. (2023, 2024). Method 4 (CI+LCCSDs) starts from LCCSD-obtained potentials, keeps two-body potentials constant and applies a rescaling parameter to the one-body potential. Method 5 (CI+CCSDs) does the same but starts from CCSD-obtained potentials. The rescaling parameters $\rho_{\kappa}$ for CI+LCCSDs and CI+CCSDs are listed in Table 2. Although CI+LCCSDs and CI+CCSDs methods depend on experimental energies, they remain valuable tools. They can provide high-lying excitation energies with appreciable accuracy, and by comparing results obtained with and without scaling parameters we can see how sensitive the property is to higher-order correlation, yielding a rough uncertainty estimate. Since the present work targets some low-lying states, we mainly use these two methods to assess the uncertainty of hyperfine-interaction parameters.
@@ III.1 Energy
From Table 3, for the Mg atom, CI+MBPT, CI+LCCSD, and CI+CCSD all yield relatively accurate results, with differences from measured values less than 200 cm-1. CI+LCCSD has the smallest difference, suggesting a cancellation between unconsidered correlation effects in CI+CCSD and non-linear terms of single and double excitations. As atomic number $Z$ increases for the three atoms, the difference between CI+MBPT results and measured values rises from 200 cm-1 to 2500 cm-1. The difference for CI+LCCSD results also increases but less significantly. CI+CCSD results are relatively stable, with differences within 250 cm-1, possibly due to more complete correlation effects. CI+LCCSDs and CI+CCSDs, the two methods with scaling parameters, have even smaller differences (controlled within 100 cm-1) from measured values.
@@ III.2 Magnetic dipole hyperfine-structure constant
$3s3p$ $^{3}\!P^{o}_{1}$ –143(1) –144.945(5) Lurio (1962) –141.365 Veseth (1987),–146.1 Porsev and Derevianko (2004)
$3s3p$ $^{3}\!P^{o}_{2}$ –127(1) –128.440(5) Lurio (1962) –126.287 Veseth (1987),–129.7 Porsev and Derevianko (2004)
@@
The last column of Table 5 shows other theoretical results. For the $3s4s$ $^{3}\!S_{1}$ state of 25Mg, there’s one available result from CI+MBPT calculation He et al. (2009). Their method only considered RPA, while ours includes RPA, HO, and TP corrections. Considering only RPA, our CI+LCCSD result (323.8 MHz) differs from theirs (325 MHz) by about 1 MHz. For $3s3p$ $^{3}\!P^{o}_{1,2}$ states of 25Mg, Veseth’s many-body perturbation theory results Veseth (1987) differ from ours by $1\%-2\%$. The CI+MBPT results by Beloy et al. Beloy et al. (2008) are consistent with ours, with a difference of less than $1$ MHz. The CI+MBPT results by Porsev et al. Porsev and Derevianko (2004) also match our recommended values.
For the 87Sr atom, our final recommended values are very close to the MCDF results of Lu et al. Lu et al. (2019), and the CI+MBPT results of Poserv et al. Porsev and Derevianko (2004), with a difference of less than $1$ MHz. The CI+MBPT results of Beloy et al. Beloy et al. (2008) are very close to our CI+MBPT results. We have not found that the hyperfine parameters of the $5s4d$ $^{1}\!D_{2}$ state have been calculated by other theoretical methods. In contrast, in the 137Ba atom, the agreement between the CI+MBPT results of Kozlov et al.  Kozlov and Porsev (1999) and the experimental results is not as good as that of our CI+LCCSD results. This indicates that these states of the 137Ba atom are more sensitive to electron correlation effects, and it also shows that the electron correlation effects beyond the second-order many-body perturbation theory are important. In general, the configuration interaction plus coupled-cluster method can accurately calculate the hyperfine-structure properties of alkaline-earth atom systems.
@@ III.3 Nuclear quadrupole moment
Table 6: The electric-field gradients $q$ (in a.u.) of the low-lying states in 25Mg , 87Sr, and 137Ba. Multiply the electric-field gradient values by 234.9648867 to obtain MHz.
@@
Table 6 shows electric-field gradients $q$ (in a.u.) of low-lying states of Mg, Sr, and Ba atoms. To express the electric-field-gradient values in MHz, multiply them by 234.9648867. Like Table 5, it lists five methods’ calculation results, final recommended values, and uncertainties. Like the hfs constant $A$, we adopt the CI+LCCSD result as our final recommended value. We compare it with the results from the other three methods and take the maximum difference as the uncertainty. For Mg, CI+MBPT method’s results differ from the other four by about 3.8%, larger than for hfs constants $A$; differences among the other four are within 1%. In Sr’s $5s5p$ $^{3}P^{o}_{1}$ and $5s5p$ $^{3}P^{o}_{2}$ states, CI+MBPT results differ from the other four by about 10%, while the other four differ within 2%, similar to the hfs constant $A$ results in Table 4. For Sr’s $5s4d$ $^{1}\!D_{2}$ state, the electric-field gradient’s electron-correlation trend differs from the hfs constant $A$. CI+MBPT results differ from the other four by about 4%, but for hfs constant $A$, CI+MBPT result is four times smaller. In Ba’s $6s5d$ $^{3}\!D_{1}$, $6s5d$ $^{3}\!D_{2}$, and $6s5d$ $^{3}\!D_{3}$ states, the five methods’ results differ within 3%, and CI+MBPT result differs from the final recommended value within 2%. For Ba’s $6s5d$ $^{1}\!D_{2}$ and $6s6p$ $^{1}\!P^{o}_{1}$ singlet states, CI+MBPT results differ from the final recommended values by 8.5% and 25% respectively, showing importance of correlation effects beyond MBPT(2). The other four methods’ results differ within 3%. Comparing Table 4 and Table 6, there are differences in electron correlation trends between magnetic-dipole and electric-quadrupole hyperfine interactions.
@@
By combining the electric-field gradients in Table 6 and the measured electric quadrupole hyperfine-structure constants of these low-lying states Lurio (1962); Grundevik et al. (1983); Zu Putlitz (1963); Heider and Brink (1977); Gustavsson et al. (1979); Schmelling (1974); Wijngaarden and Li (1995), we obtain the nuclear quadrupole moments of 25Mg, 87Sr, 135Ba and 137Ba, which are listed in Table 7. In Table 7, we not only present the nuclear quadrupole moments extracted from individual states, but also provide the final recommended values and their uncertainties. The final recommended values and their uncertainties are obtained by the following equations:
$\begin{cases}Q_{\rm Final}=\frac{\sum_{i}{\frac{Q_{i}}{(\Delta{Q_{i}})^{2}}}}{\sum_{i}{\frac{1}{(\Delta{Q_{i}})^{2}}}}\\ \Delta{Q_{\rm Final}}=\frac{1}{\sqrt{\sum_{i}{\frac{1}{(\Delta{Q_{i}})^{2}}}}}\end{cases},$ (21)
where $Q_{i}$ and $\Delta{Q_{i}}$ denote the nuclear quadrupole moment and its uncertainty obtained from the $i$-th state, respectively. The accuracy of these measured hfs constants $B$ is high. Therefore, the uncertainties of the nuclear quadrupole moments we extracted are entirely due to the electric-field gradients.
For 25Mg, the nuclear quadrupole moments extracted from the $3s3p$ $^{3}\!P^{o}_{1}$ and $3s3p$ $^{3}\!P^{o}_{2}$ states are very close to the final recommended value. The differences among the three are less than 0.2%, and the uncertainty of the final recommended value is approximately 1%. For 87Sr, the differences among the nuclear quadrupole moments obtained from $5s4d~^{1}\!D_{2}$ and $5s5p~^{3}\!P_{1,2}$ states are within 1%.
@@
The comparison shows that our result is in excellent agreement with the nuclear quadrupole moments extracted from the $6p_{3/2}$ state of 137Ba+ Silverans et al. (1986); Wendt et al. (1988); Sahoo (2006b). However, there is a deviation of about 4% compared with the results extracted from the $5d_{3/2,5/2}$ states of 137Ba+ Sahoo et al. (2013); Chakraborty et al. (2026). For the $6p_{3/2}$ state, two groups independently carried out the measurement of its electric quadrupole hyperfine-structure constants Wendt et al. (1988); Villemoes et al. (1993). The measurement results they obtained are in very good agreement, and the measurement precision of both groups is better than 1%. Similarly, for the $5d_{3/2,5/2}$ states, two teams independently measured the hyperfine splittings of these two states Silverans et al. (1986); Lewty et al. (2013). The values of the hfs constants $B$ they obtained are highly consistent, and the measurement precision of the $5d_{3/2,5/2}$ states is at least an order of magnitude higher than that of the $6p_{3/2}$ state. If the theoretical framework can accurately provide the electric-field gradient, then by combining the hfs constants $B$ of the $6p_{3/2}$ or $5d_{3/2,5/2}$ states, the nuclear quadrupole moment can be extracted accurately. In the recent work by Chakraborty et al. Chakraborty et al. (2026), they used the coupled-cluster singles, doubles, and triples method to calculate the electric-field gradient of the $5d_{3/2,5/2}$ states of 137Ba+, and obtained an accurate nuclear quadrupole moment of 137Ba. The uncertainty of their result is extremely small. Our result is obtained by statistically averaging the results of five states of the 137Ba atom. Except for the $6s5d$ $^{1}\!D_{2}$ state, the difference between our calculated magnetic-dipole hyperfine-structure constant and the measured value is within 2% as shown in Table 5. Based on this, we believe that our result should be reliable. However, there is still room for improvement in our method. For example, the contribution of the triple-excitation is not currently included in our CC calculation, and the three-body interaction potential is also ignored, which may be one of the reasons for the 4% difference. The recent work by Chakraborty et al. shows that triple excitations account for $1-2\%$ of the hfs constants $B$ of the $5d_{3/2}$ and $5d_{5/2}$ stetes in Ba+ Chakraborty et al. (2026). A contribution of this size is expected to be equally relevant for neutral alkaline-earth atoms. Since the situation of 135Ba is the same as that of 137Ba, we only compared 137Ba in Table 8.